A drawer holds 4 red hats and 4 blue hats. What is the probability of getting exactly three red hats or exactly three blue hats when taking out 4 hats randomly out of the drawer and immediately returning every hat to the drawer before taking out the next?
Answer(s): C
Getting three red out of 4 that are taken out has 4 options (4! / (3!*1!)) each option has a probability of (1/2)4 since drawing a red or blue has a 50% chance. 4*1/16= 1/4 to get three red hats. The same goes for three blue hats so 1/4+1/4 =1/2. The correct answer is C.
If 4XZ + YW = 3 and XW + YZ = 6, what is the value of the expression (2X + Y)(2Z + W)?
(2X + Y) (2Z + W) = 4XZ + 2XW + 2ZY + WY.Now, plug in this data to get: 3 + 2 x 6 = 15. The correct answer is C.
X is an even number and Y is a positive odd number. Which of the following expressions cannot be even?
Answer(s): E
The fastest way to solve this problem is by plugging in some numbers. Let’s say: X = 2, Y = 1.According to answer e: 1 x 1 = 1 and that must be an odd number. The correct answer is E.
How much interest will $2,400 earn at an annual rate of 8% in one year if the interest is compounded every 4 months?
Here, it is enough to calculate the simple interest of 8% per year.$2,400 x 8/100 x 1 = $192. Since we are calculating as a compounded rate, the interest should be a bit higher or C as the correct answer.
X is a prime number bigger than 10. Also, Y = X+X3+X5+X7. What is definitely true about Y?
Because X is a prime number bigger than 10, it must be odd. Ignoring the powers of X in the expression of Y, you’ll see that Y is a sum of 4 odd numbers therefore it must be even. The correct answer is C.
Share your comments for GMAT GMAT SECTION 2: QUANTITATIVE exam with other users:
wonderful questions
i used this site since 2000, still great to support my career
why is the answer to "which of the following is required by scrum?" all of the following stated below since most of them are not mandatory? sprint retrospective. members must be stand up at the daily scrum. sprint burndown chart. release planning.
great job. hope this helps out.
upload please. many thanks!
this is so interesting
great material thanks
anyone who wrote this exam recently
ok they re good
relevant questions
please post
q:42 there has to be a image in the question to choose what does it mean from the options
looking for cphq dumps, where can i find these for free? please and thank you.
@aarun , thanks for the information. it would be great help if you share your email
1z0-1078-23 need this dumps
i gave the microsoft azure az-500 tests and prepared from this site as it has latest mock tests available which helped me evaluate my performance and score 919/1000
i cannot see the button to go to the questions
good questions
q-6 ans-b correct. https://docs.paloaltonetworks.com/pan-os/9-1/pan-os-cli-quick-start/use-the-cli/commit-configuration-changes
very nice very nice
please help us with 1z0-1107-2 dumps
please upload the practice questions
need this dumps
preparing for this exam is overwhelming. you cannot pass without the help of these exam dumps.
new to this site but i feel it is good
the correct answer to q8 is b. explanation since the mule app has a dependency, it is necessary to include project modules and dependencies to make sure the app will run successfully on the runtime on any other machine. source code of the component that the mule app is dependent of does not need to be included in the exported jar file, because the source code is not being used while executing an app. compiled code is being used instead.
Delayed the exam until December 29th.
A and D are True
good one with explanation
This is one of the most useful study guides I have ever used.